OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 6) and (1, 3)?

OpenStudy (nikato):

First u want to find the slope

OpenStudy (nikato):

If u don't know, check this out http://www.purplemath.com/modules/slope.htm

OpenStudy (the_fizicx99):

Do you know the formula? Or is there something you don't understand on how to go about it

OpenStudy (anonymous):

(y - 3) = -3(x - 1) (y + 3) = 3(x + 1) (y + 3) = -3(x + 1) (y - 3) = 3(x + 6)

OpenStudy (nikato):

So the first thing u want to do is find the slope of those 2 points

OpenStudy (the_fizicx99):

Plug in your coordinates, M = Y2-Y1/X2-X1 to find the slope then plug in for y-y1=m(x-x1)

OpenStudy (anonymous):

whats the answer

OpenStudy (the_fizicx99):

I can't give it to you.